NH3 converts into N2 and H2 according to the following equation:-2NH3 = N2 + 3H2x% of NH3 is converted into product. Rate if diffusion of gaseous mixture is double than that of SO2 under identical conditions. Calculate x. (Ans- 6.25%)

Grade:12th pass

1 Answers

Raghvendra Kumar Suman
3 years ago
We know acc. To Graham's law of diffusion.
R1/R2=(M2/M1)^1/2
           2NH3       ->    N+   3H2
At t=0    a
At t=t     a-ax                  ax/2     3ax/2
Molar mass of mixture, 
Mmix. = (w1+w2+w3)/(n1+n2+n3)
           =[ax/2*28+ 3ax/2+a (1-x)17]/a (1+x)
           =17/(1+x)
  Rmix/R SO2  = (MSO2/ Mmix)^1/2
  2R/R= [64/(17/1+x)]^1/2
          》》x=0.0625 or 62.5%